

********************************************************************
UARS WEEKLY STATUS REPORT - 28 JAN 2005
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to perform nominally, taking solar data on all
available orbital tracking periods. This week we received about 65% of
each day's data amounting to nearly 486 minutes of solar observing time
per day.
***********
SUSIM
***********
SUSIM successfully gathered daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring scans
for time periods with UARS telemetry or tape recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004259 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004260 and 2004269 are waiting archival at GSFC DAAC.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004239. MEPS data between 2004240 and 2004259
are waiting for archival at the DAAC. MEPS data between 2004260 and 2004269
are currently being processed.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 15 January 2005 (Orbit 73045, GMT Day 015) through
21 January 2005 (Orbit 73150, GMT Day 021).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SOLSTICE 015/1406 73054 DAY ONLY TO FULLTIME
SUSIM 015/1548 73055 POWER ON
HALOE 017/1259 73083 POWER ON
The beta angle ranged between the angles of 12.2 degrees to 30.7
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 39 degrees on 27 January 2005 (DOY 027).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.65 to 1.00 degrees C. Additional
battery performance data for each battery on 14 January 2005 and 21
January 2005 are provided in Appendix D.
The clock error ranged from 10.1 to -10.1 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 03
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 09
SOLSTICE = 01
SUSIM = 02
PEM = 01
HALOE = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 05
Three (3) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to closely monitor the batteries. On
Monday afternoon the Battery-2 temperature broke 8 degrees for the
second consecutive orbit, so the MPS was commanded back to VT-6. The
battery differentials were nominal for the first several orbits after
the change and then began to push the -300 mV level by the 018/1724 Z
event. The FOT met with the NASA power branch engineer and decided
the maximum differential allowable before going back to VT-7. That
limit was exceeded on the 019/0852 event when the differential voltage
on Battery-3 hit -459 mV. The MPS was then commanded back to the VT-7
charge rate. Since that time, the batteries have performed nominally
and the MPS continues to utilize the VT-7 charge rate.
The FOT also continued their CMS software acceptance testing this
week. The goal for completion is no later than end of January 2005.
MEETINGS
The FOT met with the NASA power branch engineers on 19 January 2005 to
discuss the short and long term plans to manage the UARS batteries.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 12 February 2005 (DOY 043).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 21 January 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
73054 015/1406 05-017 SOLSTICE DAY ONLY TO FULLTIME OPS
73054 015/1406 05-015 PMON ACTION #23 RESET/ENABLE
73055 015/1548 05-018 SUSIM POWER ON - PART-1 START
73055 015/1627 05-018 SUSIM POWER ON - PART-1 COMPLETED
73060 015/2320 05-019 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
73068 016/1223 05-020 PEM POWER ON - DAY ONLY
73068 016/1303 05-018 SUSIM POWER ON - PART TWO
73068 016/1313 05-021 SUSIM POWER ON - SUPPLEMENT (START)
73069 016/1337 05-021 SUSIM POWER ON - SUPPLEMENT (FINISH)
73079 017/0534 05-022 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
73083 017/1259 05-023 HALOE POWER ON - PART 1
73086 017/1726 05-024 HALOE POWER ON - PART 2
73087 017/1948 05-025 MPS VT L-7 TO VT L-6
73088 017/2057 05-026 MPS BATSOC
73111 019/0852 05-027 MPS VT LEVEL 6 TO VT-7
73119 019/2231 05-028 MPS PMON ACTION CLEAR
73135 021/0007 05-029 MACS RE-ENABLE 4 WHEEL CONTROL
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 21 January 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
OBSPMACENA23 PMON ACTION 23 OOL 73100
OBSPMGPOOL31 PMON GROUP 31 OOL 73100
OBSRWY PITCH WHEEL FAIL 73134
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 21 January 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
73060 015/2328 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
73079 017/0534 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 21 January 2005
14 January 2005 - GMT Day 014 Beta = 08.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 67.2/ 00.0 +6.43/+4.82 +28.0/-12.8 26.4 1.03 18.9
3 -33.6/-313.6 +4.50/+2.91 +26.0/-11.2 26.4 1.04 17.5
21 January 2005 - GMT Day 021 Beta = 30.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-1.80 +00.0/+00.0 07.0 0.00 0.0
2 50.4/ -11.2 +9.39/+8.39 +28.4/-15.2 26.7 1.05 21.0
3 -00.0/-123.6 +7.08/+6.43 +26.4/-14.8 26.7 1.04 19.7
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
73087 017/1948 LV7 -> LV6
73111 019/0852 LV6 -> LV7
APPENDIX E
UARS ANOMALY REPORTS
Week ending 21 January 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-001 FOT 73056 015/1807 30832 Generic Late Acquisition
05-001 FOT 73094 018/0525 30832 Generic Late Acquisition
05-003 MPS 73100 018/1538 N/A PMON GRP 31 AND ACTION 23 OOL
05-001 FOT 73124 020/0628 30832 Generic Late Acquisition
05-004 MACS 73134 020/2122 N/A Pitch Wheel Anomaly
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-001 FOT See List Jan 05 30832 Generic Late Acquisition
05-002 MPS 73008 012/1251 NA BATT 3 Diff Voltage High
05-003 MPS 73100 018/1538 N/A PMON GRP 31 AND ACTION 23 OOL
05-004 MACS 73134 020/2122 N/A Pitch Wheel Anomaly